By itself the cold water faucet can fill a tub in 9 minutes, while by itself the hot water faucet can fill the same tub in 7 minutes. together the two faucets can fill the tub in ____ minutes
Answer by venugopalramana(3286)

By itself the cold water faucet can fill a tub in 9 minutes
CWF......IN ....1 MT.....FILLS 1/9 TUB
, while by itself the hot water faucet can fill the same tub in 7 minutes.
HWF.....IN 1 MT....FILLS.......1/7 TUB
TOGETHER THEY FILL IN 1 MT....1/9 +1/7 = (9+7)/63=16/63 TUB
TOGETHER THEY NEED 63/16 MTS = 3.94 MTS....TO FILL 1 TUB
together the two faucets can fill the tub in __3.94__ minutes